OpenStudy (anonymous):

use addition method. 5x+3y=2 3x+5y=-2

HanAkoSolo (jamierox4ev3r):

just add the two together then... 5x+3x=8x (hint) :)

HanAkoSolo (jamierox4ev3r):

@ashleysmith

HanAkoSolo (jamierox4ev3r):

lol can u figure out based on this? @ashleysmith ?

OpenStudy (anonymous):

i got 8x+8y=0

OpenStudy (anonymous):

how do i get x and y

HanAkoSolo (jamierox4ev3r):

well if 8x and 8y equal zero, there is an automatic assumption that both variables equal zero as they do not produce a value other than zero... but you check by solving for each variable.

HanAkoSolo (jamierox4ev3r):

8x+8y=0 -8x -8x -------- 8y=-8x+0 /8 /8 /8 -------- y=-x+8 y=0 because solving for x would yield the same thing, so x also equals y. Does that make sense @ashleysmith ?

OpenStudy (anonymous):

yeah thank you! @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

lol no problem :)